OpenStudy (anonymous):

Limits question:

OpenStudy (anonymous):

\[1/\sqrt{x} \]

OpenStudy (anonymous):

point (1,1)

OpenStudy (ipwnbunnies):

Limit as x approaches 1? o.o

OpenStudy (anonymous):

yup.

OpenStudy (ipwnbunnies):

Direct substitution works fine here. You even gave your own answer. Since there exists a point at that x-value, that'll also be the limit of the function at x=1.

OpenStudy (anonymous):

But, the answer is suppose to be -1/2

OpenStudy (ipwnbunnies):

No it's not. :3 http://www.wolframalpha.com/input/?i=limit+as+x+approaches+1+of+1%2Fsqrt%28x%29

OpenStudy (anonymous):

HMMMMMMMMMM. alright, thanks. lol

OpenStudy (ipwnbunnies):

Where'd you get -1/2 from? lol

OpenStudy (anonymous):

From my textbook. The answers at the back.

OpenStudy (ipwnbunnies):

The lies are real. Interesting. Are you sure you wrote it right, or you're looking at the wrong question/answer?

OpenStudy (anonymous):

I'm looking at the right question.

OpenStudy (ipwnbunnies):

...The wrong answer? :3 Sorry, I'm not sure at all y dey lie.

OpenStudy (anonymous):

Lol, alright, thanks.

OpenStudy (anonymous):

Actualllllllllllly, we're suppose to use this equation: f(x)-f(a)/x-a. Then we're suppose to sub. in everything. and I get up until here: \[-\sqrt{x}/\sqrt{x}(x-1)\] andddddd then you're suppose to sub. in 1 to see if it is one of the odd cases, which it is. But I don't know where to go from there.

OpenStudy (dumbcow):

what are you using for "a" ?

OpenStudy (anonymous):

a would be 1.

OpenStudy (anonymous):

and then f(a) would be 1 too :P

OpenStudy (dumbcow):

oh shoot i used the wrong f(x) :{

OpenStudy (anonymous):

oops. :P

OpenStudy (dumbcow):

ok you should have \[\frac{1-\sqrt{x}}{\sqrt{x}(x-1)}*\frac{(1+\sqrt{x})}{(1+\sqrt{x})} = \frac{-(x-1)}{\sqrt{x}(x-1)(1+\sqrt{x})}\]

OpenStudy (anonymous):

yeah I did. now.

OpenStudy (anonymous):

do*

OpenStudy (anonymous):

what would my restriction be for \[1+\sqrt{x}\]

OpenStudy (dumbcow):

no restriction, you are taking limit at x=1

OpenStudy (anonymous):

But, i always have to write what the denominator cannot be so it doesn't equal 0.

OpenStudy (dumbcow):

oh i see well "1 + sqrtx" is never 0 so its not an issue here

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Merci beaucoup!